OpenStudy (anonymous):

Which two points satisfy y = -x2 + 2x + 4 and x + y = 4?

OpenStudy (welshfella):

solve by substitution from the second equation y = 4 - x therefore 4 - x = - x^2 + 2x + 4 solve this for x and then find y using this result

OpenStudy (welshfella):

ok?